# A Survey of Route Recommendations: Methods, Applications, and Opportunities

### 01 Mar 2024 v1

@@ 4.4 Graph-based Approaches
Besides temporal correlations in route-related data, spatial information is also a critical factor in route recommendation tasks. In urban situations, many data have strong structural characteristics and can be naturally modeled as temporal-spatial graphs, such as knowledge graphs and relational recommendations. Many urban transportation networks are irregular, large-scale, and expanding, which fail many traditional machine learning models. In response to such new challenges, graph neural networks (GNNs) (122) recently have shown great potential in temporal-spatial research. GNNs are used to process graph-structured data which can broadly include numerous data types such as networks, texts, molecular biology, chemistry, and physics. In urban computing scenarios, routes, trajectories, and road information can be naturally modeled by GNNs with or without additional neural network modules (e.g. RNNs, CNNs, deep encoders, or attention-based modules). This subsection presents GNN-based work for route recommendation. The general process is shown in Figure 5.
@@
Many studies have applied GNNs to urban route recommendation tasks, which have considered various factors such as road traffic information, user preferences, historical trajectories, etc. For example, Wang et al. (74) used graph convolutional networks (GCNs) to represent traffic information and extract user preferences from historical trajectories for personalized route recommendations. Gao et al. (92) used graph a spatio-temporal network to fuse POI and track location information, and employed transfer learning techniques to address the data sparsity problem in travel route recommendation. Wang et al. (75; 81) used a GNN to estimate the road distance, travel time, and heuristic function $h(x)$ in A* based on the extracted features of the GNN. Wen et al. (82) employed a GCN to encode the checkpoints and trajectories along goods delivery paths and extracted the spatio-temporal dependencies for providing more profitable routes. Wu et al. (93) addressed the traveling salesman problem by using graph attention networks for path node embedding and applied their model for delivery route recommendations. Wu et al. (94) utilized a graph attention network to extract features from road vehicle trajectory and speed information and improved upon the A* constraint method. Wu et al. (76) constructed a three-level graph neural network structure to learn the representation of different functional areas in cities and empirically verified the effectiveness of the additional features used in route recommendations. Jain et al. (78) developed NeuroMLR, which learns a generative model from historical trajectories by conditioning on three explanatory factors: the current location, the destination, and real-time traffic conditions. The conditional distributions are learned through a combination of Lipschitz embedding with a GCN using historical trajectory data. Yang et al. (79) applied graph attention networks and pre-trained models to embed graph-structured data and combined A* to compute approximate graph edit distance in a more effective and intelligent way.
@@ 4.5 Multi-modal Approaches
Most of the existing route recommendation methods process two types of data - trajectories and road network information. However, with the development of urban information systems, many more types of data are recorded, and thus single-modal methods may not be sufficient to meet the travel needs of users. Meanwhile, multi-modal learning paradigms have been shown very effective in other research areas such as computer vision (123) and natural language processing (124; 125). Therefore, multi-modal route recommendation methods are a promising research area. New modalities in urban computing include but are not limited to weather information, different types of transportation networks, scenic images, texts, and audio. In this sub-section, we review the literature from two perspectives: one mainly fuses images, texts, audio, and weather data (Figure 6), and the other fuses different types of transportation data (Figure 7).
@@
In reality, people’s travel plans can be chosen from different modes of transportation or a mixture. So, recommendation algorithms in certain scenarios need to consider such multiple choices and find the best way (or a mix) of transportation. For instance, Georgakis et al. (105) proposed a multi-modal path planner that allows different transportation modes for different parts of the travel plan. Bucher et al. (100) developed an algorithm based on user profiles and generated personalized multi-modal routes among a set of feasible travel options. Herzog et al. (102) proposed RouteMe, a mobile recommendation system for personalized multi-modal routes. They combined Collaborative Filtering (CF) with knowledge-based recommendations to enhance the quality of route recommendations. Cheng et al. (103) tackled the aggregation problem of transportation modes by modeling travel time, arrival time, bicycle ownership, and transfer time among different modes of transportation. They introduced TERP, a path planner that considers time events and optimizes both travel time and reliability. Zhang and Arentze (99) estimated the distribution of people’s different travel patterns and modes. Based on this, they recommended personalized travel choices to users that are similar to each other. Yu and Lu (98) developed an approach that automatically constructs a pedestrian network after analyzing both the pedestrian infrastructure and semantic analysis. Liu et al. (104) extracted multi-modal transportation maps from large-scale map query data. Based on this, they inferred user correlations and origin-destination correlations, and used such information to recommend online multi-modal routes. In their subsequent work (106; 107; 108; 109), they further incorporated a pre-training mechanism and a contrastive learning approach to analyze trajectories and road networks.
4.6 Reinforcement Learning Related Approaches
Reinforcement Learning (RL) is a learning paradigm that involves an agent interacting within an environment strategically so as to maximize the rewards it can receive. The agent perceives the environment’s states and acts accordingly. So the key of RL is to learn a good policy, or equivalently, a Q function $Q(s,a)$ for the agent. Modern deep RL (DRL) work approximates the Q function using deep networks, so the name. RL methodology can be naturally applied to route recommendation as it is intrinsically a decision process. For example, agents are users, states are the environment’s states, actions are different moving options, and rewards can be linked to the original objective functions. The general elements of RL for route recommendation are shown in  Figure 8.
Figure 8: Reinforcement learning Approaches.
@@ 4.6 Reinforcement Learning Related Approaches
Many RL methods are applied to urban route recommendations, which have considered various factors such as cost, distance, evacuation, driving preferences, and spatio-temporal patterns. For example, Xia et al. (110) studied the classic constrained shortest paths (CSP) problem and developed an iterative search process. They designed an adaptive controller based on RL to effectively control the expansions of candidate options in each search iteration. Chen et al. (111) proposed a generic bi-criteria optimal path-finding framework (cuRL) based on DRL, which co-optimizes two objective functions - minimized distance and optimized extra cost incurred when satisfying user’s preferences. Bi et al. (112) proposed an evacuation route recommendation method, which first uses an autoencoder to perform dimensional reduction and then employs an RL-based route-selection algorithm to design routes. Liu et al. (13) developed a deep inverse RL (IRL) algorithm to learn the driving preferences of transportation personnel from their GPS trajectories and then recommend their preferred routes. Liu and Jiang (113) further improved the IRL method for taxi routing problems by considering real-time traffic conditions. Agostinelli et al. (80) adopted RL to learn admissible heuristics in the A* search to solve path-finding problems that have large state spaces. Wu et al. (93) proposed a DRL framework to automatically learn search heuristics, which are frequently handcrafted, for routing problems. They applied their learned heuristics to the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP), shown with superior performance. Ji et al. (114) modeled a taxi routing problem as a decision process and designed an adaptive DRL method, which can better integrate spatio-temporal features and recommend more effective routes.
@@ 5.1 Tourism Route Recommendation
The key to tourism route recommendation is to match user’s interests with different attractions on a potential route. Also, common basic factors such as travel time and cost need to be considered as well. A large body of works focused on user interests (95; 126; 33; 127; 129; 47; 77; 130; 131). They typically start with extracting user’s travel interests from their historical trajectories and then recommend routes of which the POIs can best match user’s interests. The matching process uses some similarity or consistency metrics. For example, Park and Liu (95) applied the latent Dirichlet allocation (LDA) and Word2Vec models to match user interests with scenic spots. Gao et al. (126) designed an adversarial network consisting of a generator and a reviewer (i.e. discriminator) and used the reviewer to distinguish between a route representation and a user’s query representation. Huang et al. (77) constructed a heterogeneous network based on the relationship between users and POIs, and used a network embedding method to learn the features of users and POIs. Zheng et al. (51) proposed a variable neighborhood search algorithm with a hybrid particle swarm genetic optimization algorithm to recommend Top-K travel routes. Some other works still considered travel time as the main factor of routes (52; 30; 39); and some aimed to design navigation systems for traveling routes (129; 132).
@@ 5.2 Economic Route Recommendation
On taxi revenues, many works focused on how to construct an objective function that maximizes driver’s income and design different optimization strategies (62; 48; 49; 114; 68; 71; 64; 55; 73). For example, Qu et al. (62) developed a cost-effective recommending system for economic driving routes. Ge et al. (133) designed a mobile recommendation system that can recommend a range of parking locations for taxi drivers. Papadopoulos et al. (136) proposed a coordinated pricing and routing scheme for a truck routing problem, which not only considers truck drivers’ revenues but also improves the overall traffic condition in the network.
In contrast to revenue controls, many works directly study how to reduce fuel consumption  (53; 133; 56; 73). For instance, Li et al. (56) developed a Physics-guided Energy Consumption model for solving the energy-efficient path selection problem which aims to find the path with the least expected energy consumption. Nguyen and Jung (53) introduced a decentralized traffic path system based on an ant colony optimization algorithm. Their system was shown to significantly reduce travel time and fuel consumption compared to existing systems. In addition, due to the development of electric vehicles (EVs), route recommendation for EVS poses a unique challenge that we need to consider the dynamic relationship between the charging locations and real-time power consumption (32; 70; 134). One representative work was by Gareau et al. (32) that focused on route planning for EVs. They aimed to find the shortest path between two given locations while passing by several charging stations. The total time has considered the total driving time, the charging time, and the waiting time.
Crowdsourcing tasks also require efficient and economic distribution of routes, so many researchers have designed different solutions to meet the needs (82; 42; 13; 93; 135; 37; 38; 136). Costa et al. (42) proposed an exact (but expensive) solution and other practical heuristic solutions for the multi-task route selection problem of crowdsourcing tasks. Liu et al. (13) applied a deep IRL algorithm to capture the preferences of traveling salesmen from their historical GPS tracks and recommend preferred routes, thus improving the delivery efficiency.
Lastly, traveling collectively in a group, such as ride-sharing, can largely reduce individual costs. There are works (137; 69) that aimed to design a suitable route with multiple stop points for users traveling or commuting together. Since the goal is to minimize the total cost of individuals and also the ride-provider, they modeled this as a multi-objective optimization problem to solve.
@@ 5.3 Personalized Route Recommendation
Travel distance and time are still the primary considerations of users (138; 139; 140; 141; 74; 110; 76; 142), while other works also take into account other preferences such as fuel consumption, visual scenes, driving habits, and diversity  (113; 40; 81; 67; 85; 143). Typically, they incorporated distance, time, and preferences as constraints into their learning objectives which also fuse user-based features (extracted from trajectories) and road traffic information. For example, Zhu et al. (139) introduced the FineRoute algorithm, a personalized and time-sensitive route recommendation system. They considered three factors: user preferences, appropriate access time, and transition time in route generation, and measured the quality of a route at the appropriate travel time.
On the other hand, while most works recommend routes with similar patterns based on a user’s history, other users may prefer to discover untraveled routes. Therefore, research work based on difference (34; 35; 36) is also a branch of personalized route recommendation. Specifically, Chen et al. (34) proposed a parallel splitting and merging method (RSL-Psc), which constructs new routes by dividing and combining candidate sub-routes. The subtasks are executed in parallel to provide users with multiple different route choices.
@@ 5.4 Security-related Route Recommendation
Here we discuss two different aspects of security. One is to protect user privacy or other security issues in route recommendations (144; 145; 146). The other is to provide safe evacuation routes during emergencies. For example, Mohdnordin et al. (31) developed an ambulance routing system that aims to ensure the availability of ambulances for emergencies within a limited time. Herschelman et al. (72) proposed a new method to minimize both the evacuation time and the number of mobile conflicts. Bi et al. (112) proposed a prediction model based on the Markov decision process and aimed to design global optimal evacuation routes. Khan et al. (147) provided a scalable emergency evacuation service that can guide evacuees to a safe and least congested route.
@@ 6 Conclusions and Future Directions
Route recommendation now plays an increasingly important role in the development of intelligent transportation and smart cities. We have seen various applications being extensively researched and deployed such as public vehicle routing, traffic management, and safe evacuation. Some works are inherited from the classic methodologies (like Dijkstra’s algorithm, A*) and aim to extend them to modern situations. This line of methods has gradually evolved from using exact solutions or simple machine learning solutions to deep learning models that can deal with more complicated and larger-scale route recommendation problems. On the other hand, many more works attempt to explore new applications, cope with new problems, process new data modes, and develop new models accordingly. With richer data volumes and types, route recommendation problems can be solved on a finer-grained level, meeting more specific user needs and preferences.
@@
    Explainability. It is widely criticized that deep learning suffers from poor explainability, though it is often considered the first choice for route recommendation research. Good explainability of models and interpretable results are crucial for route designers and city managers who aren’t even machine learning practitioners (149). We think the key to mitigating this problem is to understand the mechanisms of people’s travel patterns in depth and inject such knowledge into modeling. A trend is still to combine the classic approaches (such as exact solutions and statistical analysis which enjoy good explainability and theoretical guarantees) and deep learning models which can handle more complex problems. We have seen a lot of such practices in Section 3.1, 3.2 and 4.2. Other general solutions for improving the interpretability of deep models are to do model visualization or to interpret black-box models through their output results (3).
@@
    Multi-modal learning. Though we have mentioned some multi-modal route recommendation work in Section 4.5, this direction is still under-explored. A good practice would be to fuse visual data (e.g. images, videos of POIs, traffic, travel recordings), text data (e.g. user ratings, blogs, public announcements), and audio data into standard modeling processes, given that the general multi-modal techniques are relatively mature. However, the major bottleneck lies in data collection. Often, the multi-modal data important to our modeling aren’t easy to collect because of privacy, cost, and other issues. Therefore, how to collect multi-modal data feasibly and then design new applications to meet more kinds of user needs is a promising trend.
@@
    Unified large models. Most route recommendation models only serve for specific goals, and there is a need to study unified tasks and design very general large models. Recently, affected by the amazing progress made to the large models such as GPT (150) and Midjourney (151), an interesting question would be how we can integrate them into route-designing processes such that pre-gained knowledge can be transferred and enhance current models. One example in the urban computing field is CityGPT (152) proposed in WAIC 2023, which is a large model driven by public affairs and the joint services of multiple city departments. Up until now, incorporating large models for individual uses is still an open problem. For example, an interesting service would be to provide users with a personal travel manager who fully understands the user’s needs in various situations (commuting, traveling, etc.) and can design suitable routes accordingly.

### 06 Apr 2024 v2

@@ 4.4 Graph-based Approaches
Besides temporal correlations in route-related data, spatial information is also a critical factor in route recommendation tasks. In urban situations, many data have strong structural characteristics and can be naturally modeled as temporal-spatial graphs, such as knowledge graphs and relational recommendations. Many urban transportation networks are irregular, large-scale, and expanding, which fails many classic methods models. In response to such new challenges, graph neural networks (GNNs) (129) recently have shown great potential in temporal-spatial research. GNNs are used to process graph-structured data which can broadly include numerous data types such as networks, texts, molecular biology, chemistry, and physics. In urban computing scenarios, routes, trajectories, and road information can be naturally modeled by GNNs with or without additional neural network modules (e.g. RNNs, CNNs, deep encoders, or attention-based modules). The general process is shown in Figure 10.
Currently, the main GNN types are Graph Convolutional Networks (GCNs, Equation (13)) and Graph Attention Networks (GATs, Equation (14)). Notably, GATs have employed the attention mechanism to weigh the influence of neighboring nodes, which endows GATs with enhanced flexibility and efficacy in processing various types of graphs.
$H^{(l+1)}=\sigma\left(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}H^{(l)}W^{(l)}\right)$ (13)
Where $H^{(l)}$ denotes the feature matrix of nodes at the $l$-th layer, $W^{(l)}$ the weight matrix of that layer, $\tilde{A}$ the adjacency matrix with self-connections, $\tilde{D}$ the degree matrix of $\tilde{A}$, and $\sigma$ the nonlinear activation function.
$\alpha_{ij}$ $=\frac{\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{T}\left[Wh_{i}\|Wh_{j}\right]\right)\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{T}\left[Wh_{i}\|Wh_{k}\right]\right)\right)}$ (14)
$h^{\prime}_{i}$ $=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}Wh_{j}\right),$
Where $h_{i}$ and $h_{j}$ represent the feature vectors of nodes $i$ and $j$ respectively, $W$ the shared weight matrix, $\alpha_{ij}$ the attention weight of node $i$ w.r.t. its neighbor $j$, $\mathbf{a}$ the parameter of the attention mechanism, and $\sigma$ the activation function.
@@
Many studies have applied GNNs to urban route recommendation tasks, which have considered various factors such as road traffic information, user preferences, historical trajectories, etc. For example, Wang et al. (84) used graph convolutional networks (GCNs) to represent traffic information and extract user preferences from historical trajectories for personalized route recommendations. Gao et al. (100) used graph a spatio-temporal network to fuse POI and track location information, and employed transfer learning techniques to address the data sparsity problem in travel route recommendation. Wang et al. (85; 90) used a GNN to estimate the road distance, travel time, and heuristic function $h(x)$ in A* based on the extracted features of the GNN. Wen et al. (31) employed a GCN to encode the checkpoints and trajectories along goods delivery paths and extracted the spatio-temporal dependencies for providing more profitable routes. Wu et al. (101) addressed the traveling salesman problem by using graph attention networks for path node embedding and applied their model for delivery route recommendations. Wu et al. (102) utilized a graph attention network to extract features from road vehicle trajectory and speed information and improved upon the A* constraint method. Wu et al. (86) constructed a three-level graph neural network structure to learn the representation of different functional areas in cities and empirically verified the effectiveness of the additional features used in route recommendations. Jain et al. (32) developed NeuroMLR, which learns a generative model from historical trajectories by conditioning on three explanatory factors: the current location, the destination, and real-time traffic conditions. The conditional distributions are learned through a combination of Lipschitz embedding with a GCN using historical trajectory data. Yang et al. (88) applied graph attention networks and pre-trained models to embed graph-structured data and combined A* to compute approximate graph edit distance in a more effective and intelligent way.
@@ 4.5 Multi-modal Approaches
Most of the existing route recommendation methods process two types of data - trajectories and road network information. However, with the development of urban information systems, many more types of data are recorded, and thus single-modal methods may not be sufficient to meet the travel needs of users. Meanwhile, multi-modal learning paradigms have been shown very effective in other research areas such as computer vision (130) and natural language processing (131; 132). Therefore, multi-modal route recommendation methods are a promising research area. New modalities in urban computing include but are not limited to weather information, different types of transportation networks, scenic images, texts, and audio. In this sub-section, we review the literature from two perspectives: one mainly fuses images, texts, audio, and weather data (Figure 11), and the other fuses different types of transportation data (Figure 12).
@@
In reality, people’s travel plans can be chosen from different modes of transportation or a mixture. So, recommendation algorithms in certain scenarios need to consider such multiple choices and find the best way (or a mix) of transportation. For instance, Georgakis et al. (113) proposed a multi-modal path planner that allows different transportation modes for different parts of the travel plan. Bucher et al. (108) developed an algorithm based on user profiles and generated personalized multi-modal routes among a set of feasible travel options. Herzog et al. (110) proposed RouteMe, a mobile recommendation system for personalized multi-modal routes. They combined Collaborative Filtering (CF) with knowledge-based recommendations to enhance the quality of route recommendations. Cheng et al. (111) tackled the aggregation problem of transportation modes by modeling travel time, arrival time, bicycle ownership, and transfer time among different modes of transportation. They introduced TERP, a path planner that considers time events and optimizes both travel time and reliability. Zhang and Arentze (107) estimated the distribution of people’s different travel patterns and modes. Based on this, they recommended personalized travel choices to users that are similar to each other. Yu and Lu (106) developed an approach that automatically constructs a pedestrian network after analyzing both the pedestrian infrastructure and semantic analysis. Liu et al. (112) extracted multi-modal transportation maps from large-scale map query data. Based on this, they inferred user correlations and origin-destination correlations, and used such information to recommend online multi-modal routes. In their subsequent work (114; 115; 116; 117), they further incorporated a pre-training mechanism and a contrastive learning approach to analyze trajectories and road networks.
4.6 Deep Reinforcement Learning Related Approaches
Reinforcement Learning (RL) is a learning paradigm involving an agent interacting within an environment strategically to maximize the rewards it can receive. The agent perceives the environment’s states and acts accordingly. So the key of RL is to learn a good policy, or equivalently, a Q function $Q(s,a)$ for the agent. While classic RL learns relatively simple Q functions or policies, modern RL relies on deep neural nets to handle more complex states, actions, and policies, allowing agents to make decisions in more intricate tasks and environments. Therefore, Deep Reinforcement Learning (DRL) has demonstrated more power in various research areas, including route planning. The general elements of DRL for route recommendation are shown in  Figure 13.
Figure 13: Deep Reinforcement Learning Approaches.
@@ 4.6 Deep Reinforcement Learning Related Approaches
Many DRL methods are applied to urban route recommendations, which have considered various factors such as cost, distance, evacuation, driving preferences, and spatio-temporal patterns. For example, Xia et al. (118) studied the classic constrained shortest paths (CSP) problem and developed an iterative search process. They designed an adaptive controller based on RL to effectively control the expansions of candidate options in each search iteration. Chen et al. (119) proposed a generic bi-criteria optimal path-finding framework (cuRL) based on DRL, which co-optimizes two objective functions - minimized distance and optimized extra cost incurred when satisfying user’s preferences. Bi et al. (24) proposed an evacuation route recommendation method, which first uses an autoencoder to perform dimensional reduction and then employs an RL-based route-selection algorithm to design routes. Liu et al. (13) developed a deep inverse RL (IRL) algorithm to learn the driving preferences of transportation personnel from their GPS trajectories and then recommend their preferred routes. Liu and Jiang (120) further improved the IRL method for taxi routing problems by considering real-time traffic conditions. Agostinelli et al. (89) adopted RL to learn admissible heuristics in the A* search to solve path-finding problems that have large state spaces. Wu et al. (101) proposed a DRL framework to automatically learn search heuristics, which are frequently handcrafted, for routing problems. They applied their learned heuristics to the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP), shown with superior performance. Ji et al. (121) modeled a taxi routing problem as a decision process and designed an adaptive DRL method, which can better integrate spatio-temporal features and recommend more effective routes.
@@ 5.1 Tourism Route Recommendation
The key to tourism route recommendation is to match user’s interests with different attractions on a potential route. Also, common basic factors such as travel time and cost need to be considered as well. A large body of works focused on user interests (103; 25; 45; 133; 135; 58; 87; 136; 137). They typically start with extracting user’s travel interests from their historical trajectories and then recommend routes of which the POIs can best match user’s interests. The matching process uses some similarity or consistency metrics. For example, Park and Liu (103) applied the latent Dirichlet allocation (LDA) and Word2Vec models to match user interests with scenic spots. Gao et al. (25) designed an adversarial network consisting of a generator and a reviewer (i.e. discriminator) and used the reviewer to distinguish between a route representation and a user’s query representation. Huang et al. (87) constructed a heterogeneous network based on the relationship between users and POIs, and used a network embedding method to learn the features of users and POIs. Zheng et al. (62) proposed a variable neighborhood search algorithm with a hybrid particle swarm genetic optimization algorithm to recommend Top-K travel routes. Some other works still considered travel time as the main factor of routes (63; 42; 50); and some aimed to design navigation systems for traveling routes (135; 138).
@@ 5.2 Economic Route Recommendation
On taxi revenues, many works focused on how to construct an objective function that maximizes driver’s income and design different optimization strategies (73; 59; 60; 121; 78; 81; 75; 66; 83). For example, Qu et al. (73) developed a cost-effective recommending system for economic driving routes. Ge et al. (139) designed a mobile recommendation system that can recommend a range of parking locations for taxi drivers. Papadopoulos et al. (141) proposed a coordinated pricing and routing scheme for a truck routing problem, which not only considers truck drivers’ revenues but also improves the overall traffic condition in the network.
In contrast to revenue controls, many works directly study how to reduce fuel consumption  (64; 139; 67; 83). For instance, Li et al. (67) developed a Physics-guided Energy Consumption model for solving the energy-efficient path selection problem which aims to find the path with the least expected energy consumption. Nguyen and Jung (64) introduced a decentralized traffic path system based on an ant colony optimization algorithm. Their system was shown to significantly reduce travel time and fuel consumption compared to existing systems. In addition, due to the development of electric vehicles (EVs), route recommendation for EVS poses a unique challenge that we need to consider the dynamic relationship between the charging locations and real-time power consumption (44; 80; 140). One representative work was by Gareau et al. (44) that focused on route planning for EVs. They aimed to find the shortest path between two given locations while passing by several charging stations. The total time has considered the total driving time, the charging time, and the waiting time.
Crowdsourcing tasks also require efficient and economic distribution of routes, so many researchers have designed different solutions to meet the needs (31; 53; 13; 101; 26; 48; 49; 141). Costa et al. (53) proposed an exact (but expensive) solution and other practical heuristic solutions for the multi-task route selection problem of crowdsourcing tasks. Liu et al. (13) applied a deep IRL algorithm to capture the preferences of traveling salesmen from their historical GPS tracks and recommend preferred routes, thus improving the delivery efficiency.
Lastly, traveling collectively in a group, such as ride-sharing, can largely reduce individual costs. There are works (142; 79) that aimed to design a suitable route with multiple stop points for users traveling or commuting together. Since the goal is to minimize the total cost of individuals and also the ride-provider, they modeled this as a multi-objective optimization problem to solve.
@@ 5.3 Personalized Route Recommendation
Travel distance and time are still the primary considerations of users (143; 144; 145; 146; 84; 118; 86; 147), while other works also take into account other preferences such as fuel consumption, visual scenes, driving habits, and diversity  (120; 51; 90; 77; 93; 148). Typically, they incorporated distance, time, and preferences as constraints into their learning objectives which also fuse user-based features (extracted from trajectories) and road traffic information. For example, Zhu et al. (144) introduced the FineRoute algorithm, a personalized and time-sensitive route recommendation system. They considered three factors: user preferences, appropriate access time, and transition time in route generation, and measured the quality of a route at the appropriate travel time.
On the other hand, while most works recommend routes with similar patterns based on a user’s history, other users may prefer to discover untraveled routes. Therefore, research work based on difference (22; 46; 47) is also a branch of personalized route recommendation. Specifically, Chen et al. (22) proposed a parallel splitting and merging method (RSL-Psc), which constructs new routes by dividing and combining candidate sub-routes. The subtasks are executed in parallel to provide users with multiple different route choices.
@@ 5.4 Security-related Route Recommendation
Here we discuss two different aspects of security. One is to protect user privacy or other security issues in route recommendations. In many route-based applications, user data are frequently uploaded to clouds. To protect data and user privacy, access control, encryption algorithms, and cryptographic methods are all important aspects. Therefore, (23; 149; 150) proposed different ways to protect user privacy when using their data for route analysis based on historical data. The other is to provide safe evacuation routes during emergencies. For example, Mohdnordin et al. (43) developed an ambulance routing system that aims to ensure the availability of ambulances for emergencies within a limited time. Herschelman et al. (82) proposed a new method to minimize both the evacuation time and the number of mobile conflicts. Bi et al. (24) proposed a prediction model based on the Markov decision process and aimed to design global optimal evacuation routes. Khan et al. (151) provided a scalable emergency evacuation service that can guide evacuees to a safe and least congested route.
@@ 6 Conclusions and Future Directions
Route recommendation now plays an increasingly important role in developing intelligent transportation and smart cities. We have seen various applications being extensively researched and deployed, such as public vehicle routing, traffic management, and safe evacuation. Some works are inherited from the classic methodologies (like Dijkstra’s algorithm, A*) and extend them to modern situations. This line of methods has gradually evolved from using exact solutions or simple machine learning solutions to deep learning models that can deal with more complicated and larger-scale route recommendation problems. On the other hand, many more works attempt to explore new applications, cope with new problems, process new data modes, and develop new models accordingly. With richer data volumes and types, route recommendation problems can be solved on a fine-grained level, satisfying more specific user needs and preferences.
@@
    Explainability. It is widely criticized that deep learning suffers from poor explainability, though it is often considered the first choice for route recommendation research. Good explainability of models and interpretable results are crucial for route designers and city managers who aren’t even machine learning practitioners (153). We think the key to mitigating this problem is to understand the mechanisms of people’s travel patterns in depth and inject such knowledge into modeling. A trend is still to combine the classic approaches (such as exact solutions and statistical analysis which enjoy good explainability and theoretical guarantees) and deep learning models which can handle more complex problems. We have seen a lot of such practices in Section 3.1, 3.2 and 4.2. Other general solutions for improving the interpretability of deep models are to do model visualization or to interpret black-box models through their output results (3). In addition, demonstrating the strong generalization of models through transferability is also an interesting topic (154). For route recommendation, many tasks need to deal with multi-domain or multi-modal data, which necessitates the research of model transfer and generalization ability across different domains.
@@
    Multi-modal learning. Though we have mentioned some multi-modal route recommendation work in Section 4.5, this direction is still under-explored. Multi-modal learning can help in situations where the data of one mode are insufficient, and co-learning with other modes of data can compensate for such deficiency. A good practice would be to fuse visual data (e.g. images, videos of POIs, traffic, travel recordings), text data (e.g. user ratings, blogs, public announcements), and audio data into standard modeling processes, given that the general multi-modal techniques are relatively mature. However, the major bottleneck lies in data collection. Often, the multi-modal data important to our modeling aren’t easy to collect because of privacy, cost, and other issues. Therefore, how to collect multi-modal data feasibly and then design new applications to meet more kinds of user needs is a promising trend.
@@
    Unified large models. Most route recommendation models only serve specific goals, and there is a need to study unified tasks and design very general large models. Recently, affected by the amazing progress made to the large models such as GPT (155) and Midjourney (156), an interesting question would be how we can integrate them into route-designing processes such that pre-gained knowledge can be transferred and enhance current models. One example in the urban computing field is CityGPT (157) proposed in WAIC 2023, which is a large model driven by public affairs and the joint services of multiple city departments. Up until now, incorporating large models for individual uses is still an open problem. For example, an interesting service would be to provide users with a personal travel manager who fully understands the user’s needs in various situations (commuting, traveling, etc.) and can design suitable routes accordingly. If we particularly focus on large language models (LLMs), how to inject domain knowledge into LLMs, achieving semantic collaboration and cross-domain knowledge fusion, is an exciting direction for route recommendation. A latest survey (158) has reviewed many works that use LLMs to improve recommender systems, and similarly, the methodology could be used in route recommendation.